OpenStudy (anonymous):

write an equation for a set of points in the plane such that the sum of the distances from (-7,0) and (7,0) is 20

OpenStudy (mertsj):

This is the definition of an ellipse, right?

OpenStudy (mertsj):

Oh, I see that the asker is not present.

OpenStudy (anonymous):

I think so

OpenStudy (mertsj):

So by definition, 2a = 20. You can find the center because it is half way between the focal points.

OpenStudy (anonymous):

yes that I have done

OpenStudy (anonymous):

from here is where I am lost

OpenStudy (mertsj):

What is c?

OpenStudy (anonymous):

im not sure

OpenStudy (mertsj):

What is the center?

OpenStudy (anonymous):

well 0

OpenStudy (mertsj):

The center has an x and a y coordinate.

OpenStudy (anonymous):

yes wouldn't it be (0,0)?

OpenStudy (mertsj):

It is (0,0) And c is the distance from the center to the foci so c = 7. 2a=20 so a = 10 a^2-b^2=c^2 so 100-b^2=49 and b^2= 51. So now you know the center, a^2 and b^2. You can write the equation.

OpenStudy (anonymous):

ok thanks

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

so its (x+7)^2/51 + (y-0)^2/100=1

OpenStudy (mertsj):

a^2 goes under x^2 when the major axis is horizontal.